In a Compton scattering, the incident Cu-Kα photon has a wavelength of 1.54 Å. The ratio of the maximum kinetic energy of the recoiled electron to the energy of the incident photon will be


1
Expert's answer
2021-07-07T08:39:53-0400
"K=hf\frac{\Delta \lambda}{ \lambda+\Delta \lambda}\\\frac{K}{E}=\frac{2 \lambda}{ \lambda+2 \lambda}=\frac{2 }{3}"